Bearclaws72:

Write an expression for the perimeter of the triangle shown below: A triangle is shown with side lengths labeled 0.3x plus 24, 5x, and 0.2x minus 15. 10x − 9 10x − 39 5.5x − 9 5.5x + 9

ThisGirlPretty:

@JustSaiyan

ThisGirlPretty:

@21Pilots

JustSaiyan:

Uhhh. I don't Algebra. Sorry.

Bearclaws72:

its fine

ThisGirlPretty:

@Vocaloid You guys are making my ss rise v.v

Vocaloid:

you would just add all the sides together

JustSaiyan:

@Vocaloid Do you math?

JustSaiyan:

XD I was late. :P

Vocaloid:

0.3x + 24 + 5x + 0.2x - 15 = ?

Bearclaws72:

idk

Bearclaws72:

boi add them

Bearclaws72:

Just plug them in a calculator bird

Vocaloid:

let's start with the x's 0.3x + 5x + 0.2x = ?

Bearclaws72:

=5.5x bird

Bearclaws72:

i was using the calc lol

Vocaloid:

good 24 - 15 = ?

Bearclaws72:

9

Vocaloid:

good now we put it all together to get 5.5x + 9 = your answer

Bearclaws72:

Now theres no like terms

Bearclaws72:

14.5

Bearclaws72:

ok thnx

Vocaloid:

no, you can't add 5.5x and 9 since they're not like terms

Vocaloid:

5.5x + 9 = your answer, that's it

Bearclaws72:

ok